calculate the zero point energy of a system consisting of a mass of gm connected to a fixed point by a spring which is stretched 1cm by a force 1000 dynes the particle being constrained to move only along x axis


1
Expert's answer
2022-03-30T17:43:03-0400

"k=\frac Fx,"

"E=\frac{kx^2}2=\frac{Fx}2=50~nJ."